# Impartial Games: A Challenge for Reinforcement Learning

### 18 Dec 2023 v3

@@ Abstract
While AlphaZero-style reinforcement learning (RL) algorithms excel in various board games, in this paper we show that they face challenges on impartial games where players share pieces. We present a concrete example of a game - namely the children’s game of Nim - and other impartial games that seem to be a stumbling block for AlphaZero-style and similar self-play reinforcement learning algorithms.
@@ 1 Introduction
To gain deeper insights into the challenges faced by RL agents in mastering Nim-like games, we revisited AlphaZero-style algorithms and their variations, particularly within the context of chess. Our focus was on chess positions where the resolution of straightforward parity-related problems was crucial for accurately identifying optimal moves. This intensive exploration led to an in-depth discussion and analysis of numerous theoretical and practical issues associated with impartial games and parity problems, culminating in the development and in-depth analysis of an AlphaZero-style algorithm tailored for Nim from the champion and expert perspective, two measurements of the performance of a self-play RL agent we propose in this paper.
@@ 2 Impartial games and nim
For any Nim position, it is mathematically easy to determine which player will win and which winning moves are available to a given player, if any. The value (won or lost) of a Nim position can be determined by calculating the binary digital sum of the number of counters in the heaps, i.e., the sum in binary, neglecting all carry from one digit to another Bouton (1901). Within combinatorial game theory, this is commonly referred to as the nim-sum. The complexity of this calculation was shown to be linear time and with the use of logarithmic space memory. Thus, deciding whether a given Nim position is won or lost is computationally easy.
@@
Positions in Sprout typically have nimber values 0, 1, 2 and 3 Berlekamp et al. (2003). The position in Fig. 1(a) has nimber value of 3. The Sprout position in Fig. 1(b) also has nimber value of 3, but it has been modified to become an "isolated land" that cannot interact with anything on the outside. A Sprout starting position consisting of $n$ copies of the gadget in 1(b) can mimic any Nim position reached from a starting position with $n$ heaps of $3$ counters.
@@
Figure 2: Nimplayed on a board $[3,3,3,\ldots,3]$ with $n$ heaps, is equivalent to sprout with $n$ copies of the sprout position (gadget) in (b) in the Figure (see the diagram on p599 in Berlekamp et al. (2003) for details).
@@ 5 Reinforcement Learning for nim
There exist some open-source implementations of AlphaZero-style algorithms on GitHub  [^5] https://github.com/suragnair/alpha-zero-general [^6] https://github.com/junxiaosong/AlphaZero_Gomoku and some RL libraries also offer fast implementations Lanctot et al. (2019); Liang et al. (2018). But we needed some additional functionalities that are not provided in any of them, for instance, calculating the Elo rating of the agent being trained against its ancestors, evaluating the accuracy of output of the policy network as the training progresses against the winning move derived from the nim-sum, etc. This prompted us to implement a version of AlphaZero-style self-play algorithm in PyTorch Paszke et al. (2019) for neural network training and Ray Moritz et al. (2018) for running simulations in parallel. We made considerable efforts to ensure our implementation replicates the original AlphaZero algorithms at granular details as described in Silver et al. (2017, 2018). However, necessary changes to neural network architectures tailored to the Nim were made, which will be specified in the section 5.1.
@@ 5.2 Experiment setup and results
5.2.1 Champion measurement by Elo rating
@@
There is only one of two possible outcomes of the match for player A, being either won or lost. For nim, the drawing situation does not exist. If player A won the game, its Elo rating is updated by
@@
5.2.2 Expert measurement by the accuracy of PV-NN
In addition to the champion measurement by the Elo-rating, to examine the performance of our agent from the expert perspective, we employed two accuracy measurements on PV-NN. The action probability distribution yielded by the policy network should be biased towards the moves that have a higher chance of leading to winning the game. In Nim, the winning moves can be calculated by nim-sum, according to which the accuracy of the policy network is evaluated by comparing the accuracy of the most probable moves against the winning moves. Similar policy measurement was used in Danihelka et al. (2021) as Policy Top 1 Accuracy where the top 1 refers to the move with the highest probability.

### 14 Jan 2024 v4

@@ Abstract
While AlphaZero-style reinforcement learning (RL) algorithms excel in various board games, in this paper we show that they face challenges in impartial games where players share pieces. We present a concrete example of a game - namely the children’s game of Nim - and other impartial games that seem to be a stumbling block for AlphaZero-style and similar self-play reinforcement learning algorithms.
@@ 1 Introduction
To gain deeper insights into the challenges faced by RL agents in mastering Nim-like games, we revisited AlphaZero-style algorithms and their variations, particularly within the context of chess. Our focus was on chess positions where the resolution of straightforward parity-related problems was crucial for accurately identifying optimal moves. This intensive exploration led to an in-depth discussion and analysis of numerous theoretical and practical issues associated with impartial games and parity problems, culminating in the development and in-depth analysis of an AlphaZero-style algorithm tailored for Nim from the champion and expert perspective, two measures of the performance of a self-play RL agent we propose in this paper.
@@ 2 Impartial games and Nim
For any Nim position, it is mathematically easy to determine which player will win and which winning moves are available to a given player, if any. The value (won or lost) of a Nim position can be determined by calculating the binary digital sum of the number of counters in the heaps, i.e., the sum in binary, neglecting all carry from one digit to another Bouton (1901). Within combinatorial game theory, this is commonly referred to as the nim-sum. The complexity of this calculation was shown to be linear time Fraenkel (2004) and with the use of logarithmic space memory Calabro (2006). Thus, deciding whether a given Nim position is won or lost is computationally easy.
@@
Positions in Sprout typically have nimber values 0, 1, 2 and 3 Berlekamp et al. (2003). The position in Figure 1(a) has nimber value of 3. The Sprout position in Figure 1(b) also has nimber value of 3, but it has been modified to become an "isolated land" that cannot interact with anything on the outside. A Sprout starting position consisting of $n$ copies of the gadget in Figure 1(b) can mimic any Nim position reached from a starting position with $n$ heaps of $3$ counters.
@@
Figure 2: Nim played on a board $[3,3,3,\ldots,3]$ with $n$ heaps, is equivalent to Sprout with $n$ copies of the Sprout position (gadget). (see the diagram on p599 in Berlekamp et al. (2003) for details).
@@ 5 Reinforcement Learning for Nim
There exist some open-source implementations of AlphaZero-style algorithms on GitHub  [^5] GitHub. https://github.com/suragnair/alpha-zero-general [^6] GitHub. https://github.com/junxiaosong/AlphaZero_Gomoku and some RL libraries also offer fast implementations Lanctot et al. (2019); Liang et al. (2018). But we needed some additional functionalities that are not provided in any of them, for instance, calculating the Elo rating of the agent being trained against its ancestors, evaluating the accuracy of output of the policy network as the training progresses against the winning move derived from the nim-sum, etc. This prompted us to implement a version of AlphaZero-style self-play algorithm in PyTorch Paszke et al. (2019) for neural network training and Ray Moritz et al. (2018) for running simulations in parallel. We made considerable efforts to ensure our implementation replicates the original AlphaZero algorithms at granular details as described in Silver et al. (2017, 2018). However, necessary changes to neural network architectures tailored to the Nim were made, which will be specified in the section 5.1.
@@ 5.2 Experiment setup and results
5.2.1 Champion measure by Elo rating
@@
There is only one of two possible outcomes of the match for player A, being either won or lost. For Nim, the drawing situation does not exist. If player A wins the game, its Elo rating is updated by
@@
5.2.2 Expert measure by the accuracy of PV-NN
In addition to the champion measure by the Elo-rating, to examine the performance of our agent from the expert perspective, we employed two accuracy measures on PV-NN. The action probability distribution yielded by the policy network should be biased towards the moves that have a higher chance of leading to winning the game. In Nim, the winning moves can be calculated by nim-sum, according to which the accuracy of the policy network is evaluated by comparing the accuracy of the most probable moves against the winning moves. A similar policy measure was used in Danihelka et al. (2021) as Policy Top 1 Accuracy where the top 1 refers to the move with the highest probability.